Fermi National Laboratory
 

Accelerator Update for July 27, 2007 - July 30, 2007

 

Friday July 27

The day shift began with Operations monitoring the Tevatron (TeV) store 5582 (87.86E30), the Recycler stash (118.4E10), and the Antiproton Source (Pbar) stack (23.4E10), with no Linac beam due to problems with a RF station (LRF3), with Booster, and Main Injector (MI) waiting on Linac beam, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams, and with MiniBooNE and the Meson MTest experiments also waiting on Linac beam. NuMI will remain off until after the summer shutdown.

Operations reported at 8:43 AM that Linac experts were beginning to change out the Power Amplifier (PA) for LRF3.

At 9:40 AM, two Booster RF stations (BRF10 & 12) tripped off and would not turn on remotely. Operators reset the stations and turned them on locally. This happened three more times over the next 45 minutes.

Linac experts reported at 10:32 AM that they had completed the LRF3 PA change out.

Operations reported at 11:30 AM that they had a Linac beam permit.

Operations began sending beam to Pbar, MiniBooNE, and MTest at 11:38 AM.

Recycler experts reported at 11:48 AM that they were switching from the old to the new damper system and then conducting transfer function measurements for both systems.

TeV experts began separator scans at 1:30 PM. D0 and CDF turned off their detectors during the scans.

At 1:46 PM, BRF11 tripped off. It reset.

The TeV separator scans ended at 2:12 PM. D0 and CDF turned the detectors back on.

At 2:33 PM, BRF1 & 3 tripped off. They reset.

At 2:46 PM, BRF15 tripped off for the first of three times in six minutes. It reset each time.

At 2:50 PM, Operations reported that the air compressor at TeV sector E2 was leaking. Experts began repairs.

Recycler experts reported at 2:52 PM that they had completed their damper measurements and were beginning a different study.

A MI expert began a slipstacking study at 3:36 PM.

Operations reported at 3:55 PM that the Recycler had stashed antiprotons once during the day shift.

The MI studies ended at 5:54 PM.

At 7:30 PM, CDF reported that parts of their detector had tripped off. Initially, they thought it was due to bad beam positions, but that wasn’t the case. They weren’t sure why parts tripped off, but everything turned back on okay.

A MI expert began another round of slipstacking studies at 9:57 PM.

Operations reported at midnight that the Recycler had stashed four times during the evening shift.

Saturday July 28

The midnight shift began with Operations monitoring store 5582 (39.19E30) and the stash (253.9E10), with Pbar stacking (39.19E10), and with MiniBooNE taking beam.

Operations terminated store 5582 at 5:01 AM.

At 5:33 AM, a TeV quadrupole power supply (T:QE17) showed a fuse fault. Operators replaced its regulator.

Operations put the TeV through a dry squeeze at 6:16 AM.

Operations put the TeV into shot setup at 6:35 AM.

At 6:49 AM, Operations reported that the new Recycler damper system was in alarm. Vertical emittances grew and about 3E10 of antiprotons were lost. Operations called in an expert.

At 7:36 AM BRF3 tripped off. It reset.

At 7:47 AM, a Booster vacuum pump tripped off and caused BRF7 & 8 to trip. Everything reset.

At 7:52 AM, the P150 transfer line suffered a vacuum burst. The vacuum recovered within a few minutes.

Operations reported at 7:58 AM that the Recycler had stashed twice during the midnight shift.

Operations established store 5584 at 8:27 AM with an initial luminosity of 221.45E30.

Operations established beam to MTest at 8:33 AM.

Recycler experts turned off their new damper system at 10:16 AM to perform a transfer function measurement.

At 2:30 PM, Operations reported that a Recycler 453 control card for four corrector magnet wasn’t responding. Operators replaced the card and restored its tables.

At 4:13 PM, Operations reported that the Recycler 453 card operators had replaced earlier was showing errors. Operators replaced the card again, but this time the stash was lost, which held 120E10 of antiprotons. A Recycler expert was informed and he said that changing the card was a necessary but known risk.

At 5:57 PM, Operations reported that during a stashing transfer there was a problem. When operators restarted the sequencer for the transfer, the antiprotons were in the wrong place and were lost. An expert explained that a parameter hadn’t been set properly.

MTest finished using the beam for the night at 6:09 PM.

Operations reported at 8:02 PM that the Recycler had stashed four times during the last twelve hours. Accelerator status: store 5584 was at 68.49E30, the stack was at 45.5E10, and the stash was at 53.7E10.

At 8:38 PM, Operations reported that Pbar’s AP10 transfer line had tripped off for the first of about six trips in three hours due to an increase in beam intensity. Over this period, operators lowered the intensity at times and then raised it back to normal.

At 11:54 PM, BRF10 tripped off. It reset.

Sunday July 29

The midnight shift began with Operations monitoring store 5584 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 1:51 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.

At 2:25 AM, BRF1 tripped off. It reset.

MTest resumed taking beam at 6:31 AM.

Operations reported at 8:02 AM that the Recycler had stashed five times during the last twelve hours. Accelerator status: store 5584 was at 39.81E30, the stack was at 42.4E10, and the stash was at 209.3E10.

At 10:50 AM, a Preaccelerator expert came in to tune up the I- Source. Its output had dropped. The expert worked on it for three hours without much success. There was a known overheating problem and this was the limiting the source output.

At 11:33 AM, Operations reported losing 20E10 of antiprotons from the Recycler with a similar sequencer problem that had happened earlier.

At 11:46 AM, a Linac quadrupole (Q304) had tripped off. It reset.

At 12:06 PM, LRF2 tripped off. It reset.

Operations terminated store 5584 at 12:47 PM.

TeV experts began an aperture scan study at 1 PM.

Per a Pbar expert’s suggestion, Operators changed a beam monitor in the AP10 transfer line at 2:59 PM.

Operations put the TeV through a wet squeeze around 4:10 PM.

Operations prepared D0 and CDF for accesses at 4:31 PM.

MTest reported at 6:08 PM that they were through with the beam for the night.

Both D0 and CDF were out of access by 6:28 PM.

Operations put the TeV through a dry squeeze at 7 PM.

Operations put the TeV through a wet squeeze at 7:32 PM.

Operations reported at 7:56 PM that the Recycler had stashed four times during the last twelve hours. Accelerator status: the stack was at 41.9E10, and the stash was at 329.1E10.

Operations put the TeV into shot setup at 7:59 PM.

Operations established store 5588 at 10:17 PM with an initial luminosity of 257.98E30.

At 10:48 PM, BRF18 tripped off. It reset.

Monday July 30

The midnight shift began with Operations monitoring store 5588 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

At 1:29 AM, BRF1 & 3 tripped off. They both reset.

At 4:40 AM, LRF3 tripped off. It reset.

At 6:30 AM, the Booster West Anode Power Supply tripped off. Operators investigated and found that BRF12 was the cause. Operators bypassed the station and brought BRF19 on in its place.

Operations reported at 8:03 AM that the Recycler had stashed four times during the last twelve hours.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator might allow some machine studies today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/30/2007    email Fermilab